A carbon resistor dissipates 60 W of power from a 120 V source at 20°C. How much power will be dissipated in the resistor at 120°C if connected across the same source? Assume the temperature coefficient of carbon at 20°C is -0.0005 per °


1
Expert's answer
2021-10-15T12:42:28-0400

"P=U^2/R\to R=U^2/P=120^2/60=240\ (\Omega)"


"R=R_0(1+\alpha t) \to240=R_0(1-0.0005\cdot20) \to"


"R_0=242.42\ (\Omega)"


"R_1=242.42\cdot(1-0.0005\cdot120)=227.87(\Omega)"


"P_1=120^2/227.87=63.2\ (W)" . Answer